# Targeted Data Fusion for Causal Survival Analysis Under Distribution Shift

### v1

@@ Abstract
Causal inference across multiple data sources has the potential to improve the generalizability, transportability, and replicability of scientific findings. However, data integration methods for time-to-event outcomes—common in medical contexts such as clinical trials—remain underdeveloped. Existing data fusion methods focus on binary or continuous outcomes, neglecting the distinct challenges of survival analysis, including right-censoring and the unification of discrete and continuous time frameworks. To address these gaps, we propose two novel approaches for multi-source causal survival analysis. First, considering a target site-specific causal effect, we introduce a semiparametric efficient estimator for scenarios where data-sharing is feasible. Second, we develop a federated learning framework tailored to privacy-constrained environments. This framework dynamically adjusts source site-specific contributions, downweighting biased sources and upweighting less biased ones relative to the target population. Both approaches incorporate nonparametric machine learning models to enhance robustness and efficiency, with theoretical guarantees applicable to both continuous and discrete time-to-event outcomes. We demonstrate the practical utility of our methods through extensive simulations and an application to two randomized trials of a monoclonal neutralizing antibody for HIV-1 prevention: HVTN 704/HPTN 085 (cisgender men and transgender persons in the Americas and Switzerland) and HVTN 703/HPTN 081 (women in sub-Saharan Africa). The results highlight the potential of our approaches to efficiently estimate causal effects while addressing heterogeneity across data sources and adhering to privacy and robustness constraints.
@@ 1 Introduction
Data integration promises to improve the generalizability, transportability, and replicability of scientific findings. However, leveraging data from multiple sources is challenging because of heterogeneous data structures, covariate shifts, differing treatment assignment mechanisms, and conditional outcome distribution shifts. Furthermore, privacy regulations, such as the General Data Protection Regulation (GDPR) in Europe and the Health Insurance Portability and Accountability Act (HIPAA) in the United States, often preclude pooling individual-level data. These considerations create substantial barriers to drawing valid causal inferences, underscoring the need for robust methodologies that can accommodate multi-source data integration without violating privacy regulations.
Although recent years have witnessed rapid growth in data integration methods for causal inference (Han et al., 2021, 2024; Yang & Ding, 2019; Liu et al., ; Han et al., 2023; Li et al., 2023), comprehensive reviews emphasize that most techniques have focused on binary or continuous outcomes (Colnet et al., 2024; Brantner et al., 2023; Degtiar & Rose, 2023). When time-to-event outcomes are considered, strong assumptions such as the Cox proportional hazards model (Hernán, 2010; Han, 2023) or the common conditional outcome distribution (CCOD) across sites (Lee et al., 2022; Cao et al., 2024) are often imposed. Importantly, violating the proportional hazards assumption or mis-specifying the outcome model can lead to bias. In addition, strict exchangeability assumptions may be untenable in practice if there are unknown shifts across sites in covariates and outcome-generating mechanisms. Recently, Westling et al. (2023) integrated double/debiased machine learning (Chernozhukov et al., 2018) with semiparametric survival curve estimation, enabling flexible nonparametric models (Wolock et al., 2024; Cui et al., 2023; van der Laan et al., 2007) for nuisance function estimation. However, this method was developed in the context of a single study, and it is not obvious how one could extend the method to multi-source contexts, especially with potential privacy constraints.
To address these gaps, we propose two methodological innovations for multi-source causal survival analysis. First, under settings where data sharing is feasible, we develop a semiparametric efficient estimator (Section 2.2) that adjusts for covariate shifts while ensuring both consistency and improved efficiency. Second, in more restrictive scenarios with potential CCOD violations and limited data sharing—necessitated by policies such as GDPR and HIPAA—we introduce a federated learning framework (Section 2.3). The latter uses a data-adaptive weighting strategy, upweighting less-biased sources and downweighting sources prone to bias, without requiring pooling of individual-level data.
Our federated approach distinguishes itself from existing methods in three key ways: (i) accommodating CCOD violations and covariate shifts, even when differences among sites are unknown a priori; (ii) avoiding reliance on strong parametric or semiparametric models by using flexible machine learning for nuisance estimation while maintaining parametric convergence rates; and (iii) accommodating events in both continuous and discrete time.
@@ 2.1 Notation, estimand, and assumptions
Consider $K$ studies, each of which could be randomized or observational. For each participant in a given study, we observe a covariate vector $\mathbf{X}$ representing baseline characteristics, and a binary treatment variable $A\in\{0,1\}$, where 0 denotes control and 1 denotes the active treatment. Under the Neyman-Rubin potential outcomes framework (Neyman, 1923; Rubin, 1974), we assume each participant has two potential event times, $T^{(1)}$ and $T^{(0)}$, and two potential censoring times, $C^{(1)}$ and $C^{(0)}$, under treatments $A=1$ and $A=0$, respectively. However, we can only observe either $T=T^{(A)}=AT^{(1)}+(1-A)T^{(0)}$ or $C=C^{(A)}=AC^{(1)}+(1-A)C^{(0)}$ for each participant, assuming the stable-unit treatment value assumption (SUTVA) (Rosenbaum & Rubin, 1983). We also observe a right-censored outcome $Y=\min(T,C)$ together with an event indicator $\Delta=\mathbb{I}(T\leq C)$, where $\mathbb{I}(\cdot)$ is the indicator function. We denote an arbitrary observation as $\mathcal{O}$ and the observed data of sample size $n$ from all study sites as $\{\mathcal{O}_{i}=(\mathbf{X}_{i},A_{i},Y_{i},\Delta_{i},R_{i}),i=1,\dots,n\}$, where $R$ is a site variable taking values in $\{0,1,\dots,K-1\}$, with 0 representing the target site and the remaining $K-1$ values representing source sites, $K>1$.
In this paper, we are interested in the target-site treatment-specific survival function, denoted by $\theta^{0}(t,a)=\mathbb{P}(T^{(a)}>t\mid R=0)$ for $a\in\{0,1\}$ and $t\in[0,\tau]$ for some positive $\tau<\infty$. Here, the upper subscript 0 of $\theta^{0}(t,a)$ indicates the target site $R=0$, and $\mathbb{P}$ means the probability distribution of an ideal data unit.
@@
where
$\Lambda^{k}(t\mid a,\mathbf{x})=\int_{0}^{t}\frac{N^{k}_{1}(du\mid a,\mathbf{x})}{D^{k}(u\mid a,\mathbf{x})}$
is the conditional cumulative hazard function under $R=k$, with $N^{k}_{\delta}(t\mid a,\mathbf{x})=\mathbb{P}(Y\leq t,\Delta=\delta\mid A=a,\mathbf{X}=\mathbf{x},R=k)$ the conditional cumulative incidence function of event $(\delta=1)$ or censoring $(\delta=0)$, and $D^{k}(t\mid a,\mathbf{x})=\mathbb{P}(Y\geq t\mid A=a,\mathbf{X}=\mathbf{x},R=k)$. Notice that $S^{k}$ and $\Lambda^{k}$ have a one-to-one mapping, so by estimating $\Lambda^{k}$, we can obtain $S^{k}$, and vice versa. The product integral reduces to the normal product $\prod$ under discrete time and $\exp\{-\Lambda^{k}(t\mid a,\mathbf{x})\}$ under continuous time.
Throughout, we make the following four assumptions.
@@
Assumption 2.2 (Treatment-specific independent censoring).
@@
Assumption 2.3 (Positive proportions for sites).
There exists $\eta\in(0,\infty)$ such that $\mathbb{P}(R=k)\geq 1/\eta$, for $k=0,1,\dots,K-1$.
Assumption 2.4 (Positivity of nuisance functions).
There exists an $\eta\in(0,\infty)$, such that for $\mathbb{P}$-almost all $\mathbf{x}$, $\min\cup_{k=0}^{K-1}\{\pi^{k}(a\mid\mathbf{x}),G^{k}(t\mid a,\mathbf{x})\}\geq 1/\eta$ and $\min_{k}S^{k}(t\mid a,\mathbf{x})>0$. Here, $\pi^{k}(a\mid\mathbf{x})=\mathbb{P}(A=a\mid\mathbf{X}=\mathbf{x},R=k)$ is the propensity score of the treatment $A=a$ and $G^{k}(t\mid a,\mathbf{x})=\mathbb{P}(C>t\mid A=a,\mathbf{X}=\mathbf{x},R=k)$ is the conditional survival function of the censoring time in site $k$, for $k=0,1,\dots,K-1$, $a\in\{0,1\}$ and $t\in[0,\tau]$.
Assumption 2.3 ensures that the sampling or survey process prevents any site from contributing only an extremely small proportion of participants to the entire sample (Lee et al., 2022; Cao et al., 2024).
Under the above assumptions, the target estimand can be identified as
$\theta^{0}(t,a)=\mathbb{E}\{S^{0}(t\mid a,\mathbf{X})\mid R=0\},$
for $a\in\{0,1\}$ and $t\in[0,\tau]$. Suppose we only use data from the target site $(R=0)$, the efficient influence function (EIF) of $\theta^{0}(t,a)$ can be derived in Proposition 2.5.
Proposition 2.5 (Westling et al. (2023)).
The semiparametric EIF for $\theta^{0}(t,a)$ given $t\in[0,\tau]$ and $a\in\{0,1\}$ is $\varphi^{*0}_{t,a}(\mathcal{O};S^{0},G^{0},\pi^{0})=\varphi^{*0}_{t,a}(\mathcal{O};\mathbb{P})=\varphi^{0}_{t,a}(\mathcal{O};\mathbb{P})-\theta^{0}(t,a)$, where $\varphi^{0}_{t,a}(\mathcal{O};\mathbb{P})=$
$\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\bigg{[}1-\frac{\mathbb{I}(A=a)}{\pi^{0}(a\mid\mathbf{X})}\bigg{\{}\frac{\mathbb{I}(Y\leq t,\delta=1)}{S^{0}(Y\mid a,\mathbf{X})G^{0}(Y\mid a,\mathbf{X})}$
$-\int_{0}^{t\wedge Y}\frac{\Lambda^{0}(du\mid a,\mathbf{X})}{S^{0}(u\mid a,\mathbf{X})G^{0}(u\mid a,\mathbf{X})}\bigg{\}}\bigg{]}S^{0}(t\mid a,\mathbf{X}).$
Throughout, let $\mathbb{P}_{n}[f(\mathcal{O})]=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ be the shorthand of the empirical average and $\widehat{\mathbb{P}}$ denote that the nuisance functions are replaced by their estimates. The above EIF motivates a doubly robust estimator, given as $\mathbb{P}_{n}[\widehat{\varphi}^{0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})]$. Under regularity conditions, this estimator is regular and asymptotic linear (RAL) and achieves the semiparametric variance lower bound uniformly over $t\in[0,\tau]$.
However, this result is based on a data structure that includes only the target site. To further improve efficiency by leveraging source data from the other $K-1$ sites, we present two approaches over the next two sections.
@@ 2.2 Efficient estimation under CCOD and data sharing
When target and source sites are deemed to be sufficiently similar, it may be reasonable to assume that they have the same conditional outcome distribution, as formalized in the structural assumption below.
@@
Assumption 2.6 implies that the conditional survival function $S^{k}(t\mid a,\mathbf{X})$ has the same form across sites. It also imposes no restrictions on the covariate distribution across sites, allowing for any level of covariate shift. Under Assumption 2.6, data from non-target source sites can be leveraged to improve estimation efficiency. We first provide a result under CCOD that generalizes Proposition 2.5.
Consider additional nuisance functions (each defined without conditioning on $R$): the propensity score $\bar{\pi}(a\mid\mathbf{x})=\mathbb{P}(A=a\mid\mathbf{X}=\mathbf{x})$, the global conditional survival function for the event time $\bar{S}(t\mid a,\mathbf{x})=\mathbb{P}(T>t\mid A=a,\mathbf{X}=\mathbf{x})$, the global conditional survival function for the censoring time $\bar{G}(t\mid a,\mathbf{x})=\mathbb{P}(C>t\mid A=a,\mathbf{X}=\mathbf{x})$, and the target site propensity $q^{0}(\mathbf{X})=\mathbb{P}(R=0\mid\mathbf{X})$. Similar to Assumption 2.4, we impose the following positivity assumptions for these nuisance functions.
@@
There exists an $\eta\in(0,\infty)$, such that for $\mathbb{P}$-almost all $\mathbf{x}$, $\min\{\bar{\pi}(a\mid\mathbf{x}),\bar{G}(t\mid a,\mathbf{x})\}\geq 1/\eta$ and $\bar{S}(t\mid a,\mathbf{x})>0$, for all $a\in\{0,1\}$ and $t\in[0,\tau]$.
@@
Under Assumption 2.6, the semiparametric EIF of $\theta^{0}(t,a)$ given $t\in[0,\tau]$ and $a\in\{0,1\}$ is $\varphi^{*\text{CCOD}}_{t,a}(\mathcal{O};\bar{S},\bar{G},\bar{\pi})=\varphi^{*\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})=\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})-\theta^{0}(t,a)$, where $\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})=$
$\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}-\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\bigg{\{}\frac{\mathbb{I}(Y\leq t,\Delta=1)}{\bar{S}(Y\mid a,\mathbf{X})\bar{G}(Y\mid a,\mathbf{X})}$
$-\int_{0}^{t\wedge Y}\frac{\bar{\Lambda}(du\mid a,\mathbf{X})}{\bar{S}(u\mid a,\mathbf{X})\bar{G}(u\mid a,\mathbf{X})}\bigg{\}}\bigg{]}\bar{S}(t\mid a,\mathbf{X}).$
@@
Moreover, we employ $M$-fold cross-fitting in line with the double/debiased machine learning (DML) framework (Chernozhukov et al., 2018). We allow a number of parametric and machine learning models (or an ensemble of these) from survSuperLearner (Westling et al., 2023) and SuperLearner (van der Laan et al., 2007) packages for nuisance function esimtation.
In Theorem 2.9 below, we demonstrate the (uniform) RAL property of the proposed CCOD estimator. In the following, the subscript “$m$” denotes the prediction from the $m$-th validation set over the total sample of size $n$, where the $n$ is omitted from the subscription.
Condition 2.1.
There exists $\bar{\pi}_{\infty}$, $\bar{G}_{\infty}$, $\bar{\Lambda}_{\infty}$ and $\bar{S}_{\infty}^{k}$ such that
  1. (a)
    $\max_{m}\mathbb{P}\left[\frac{1}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}-\frac{1}{\bar{\pi}_{\infty}(a\mid\mathbf{X})}\right]^{2}\to_{p}0$;
  2. (b)
    $\max_{m}\mathbb{P}\left[{\widehat{q}^{0}_{m}(\mathbf{X})}-{q^{0}_{\infty}(\mathbf{X})}\right]^{2}\to_{p}0$;
  3. (c)
    $\max_{m}\mathbb{P}\left[\sup_{u\in[0,t]}\left|\frac{1}{\widehat{\bar{G}}_{m}(u\mid a,\mathbf{X})}-\frac{1}{\bar{G}_{\infty}(u\mid a,\mathbf{X})}\right|\right]^{2}\to_{p}0$;
  4. (d)
Condition 2.2.
There exists an $\eta\in(0,\infty)$ such that for $\mathbb{P}$-almost all $\mathbf{x}$, $\widehat{\bar{\pi}}_{m}(a\mid\mathbf{x})\geq 1/\eta$, $\bar{\pi}_{\infty}(a\mid\mathbf{x})\geq 1/\eta$, $\widehat{q}^{0}_{m}(\mathbf{x})\geq 1/\eta$, $q^{0}_{\infty}(\mathbf{x})\geq 1/\eta$, $\widehat{\bar{G}}_{m}(t\mid a,\mathbf{x})\geq 1/\eta$, and $\bar{G}_{\infty}(t\mid a,\mathbf{x})\geq 1/\eta$ with probability tending to 1.
Condition 2.3.
Condition 2.4.
Define
$\bar{r}_{n,t,a,1}$ $=\max_{m}\mathbb{P}|\{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})-\bar{\pi}(a\mid\mathbf{X})\}$
×$\qquad\qquad\times\{\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})\}|,$
$\bar{r}_{n,t,a,2}$ $=\max_{m}\mathbb{P}|\{\widehat{q}^{0}_{m}(\mathbf{X})-q^{0}(\mathbf{X})\}$
×$\qquad\qquad\times\{\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})\}|,$
$\bar{r}_{n,t,a,3}$ $=\max_{m}\mathbb{P}\bigg{|}\widehat{S}^{k}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\left\{\frac{G^{k}(u\mid a,\mathbf{X})}{\widehat{G}^{k}_{m}(u\mid a,\mathbf{X})}-1\right\}$
×$\qquad\qquad\times\left(\frac{S^{k}}{\widehat{S}^{k}_{m}}-1\right)(du\mid a,\mathbf{X})\bigg{|}.$
Condition 2.5.
Theorem 2.9 (RAL of the CCOD estimator).
If Conditions 2.12.2 hold with $\bar{\pi}_{\infty}=\bar{\pi}$, $q^{0}_{\infty}=q^{0}$, $\bar{G}_{\infty}=\bar{G}$ and $\bar{S}_{\infty}=\bar{S}$, and Condition 2.4 also holds, then $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{t,a})+o_{p}(n^{-1/2})$. In particular, $n^{1/2}(\widehat{\theta}^{\text{CCOD}}_{n}(t,a)-\theta^{0}(t,a))$ convergences in distribution to a normal random variable with mean zero and variance $\sigma^{2}=\mathbb{P}[(\varphi^{*\text{CCOD}}_{t,a})^{2}]$. If Conditions 2.3 and 2.5 also hold, then

### v2

@@ Abstract
Causal inference across multiple data sources offers a promising avenue to enhance the generalizability and replicability of scientific findings. However, data integration methods for time-to-event outcomes, common in biomedical research, are underdeveloped. Existing approaches focus on binary or continuous outcomes but fail to address the unique challenges of survival analysis, such as censoring and the integration of discrete and continuous time. To bridge this gap, we propose two novel methods for estimating target site-specific causal effects in multi-source settings. First, we develop a semiparametric efficient estimator for settings where individual-level data can be shared across sites. Second, we introduce a federated learning framework designed for privacy-constrained environments, which dynamically reweights source-specific contributions to account for discrepancies with the target population. Both methods leverage flexible, nonparametric machine learning models to improve robustness and efficiency. We illustrate the utility of our approaches through simulation studies and an application to multi-site randomized trials of monoclonal neutralizing antibodies for HIV-1 prevention, conducted among cisgender men and transgender persons in the United States, Brazil, Peru, and Switzerland, as well as among women in sub-Saharan Africa. Our findings underscore the potential of these methods to enable efficient, privacy-preserving causal inference for time-to-event outcomes under distribution shift.
@@ 1 Introduction
Data fusion, or data integration, promises to improve the generalizability, transportability, and replicability of scientific findings. However, leveraging knowledge from multiple data sources is challenging because of heterogeneous data structures and distribution shift. Furthermore, privacy regulations, such as the General Data Protection Regulation (GDPR) in Europe and the Health Insurance Portability and Accountability Act (HIPAA) in the United States, often preclude pooling individual-level patient data.
Despite rapid growth in data integration methods for causal inference [25, 26, 52, 35, 27, 33], most approaches have been developed for binary or continuous outcomes [11, 5, 16], which do not capture the timing of events. In many clinical and policy-relevant settings, the timing of an outcome—such as time to hospitalization, relapse, or death—is often as important as the occurrence itself. Methods that ignore this temporal dimension risk overlooking treatment effect heterogeneity over time and can limit their utility for decision-making by patients, clinicians, and policymakers. While some recent work has extended data fusion techniques to survival outcomes, these methods impose strong parametric or structural assumptions, such as the Cox proportional hazards (PH) model [29, 23] or a common conditional outcome distribution (CCOD) across sites [32, 7]. Violations of PH or heterogeneous outcome models can result in biased estimates of treatment effects. Moreover, assumptions of strict exchangeability across sources—frequently adopted in the data fusion literature—may fail in the presence of covariate shift, heterogeneous censoring mechanisms, or other types of distribution shift.
Traditionally, time-to-event outcomes have been analyzed using data from single-site randomized clinical trials, with survival curves estimated via the unadjusted Kaplan–Meier estimator [31]. In studies with rich covariate information, extensions have included semiparametric models like the Cox PH model [14, 51, 6] and doubly robust estimators [3]. More recently, Westling et al. [48] integrated double/debiased machine learning (DML) [8] to leverage flexible nuisance estimators [50, 15, 45] to estimate survival curves. However, this approach was designed for single-study settings, and it remains unclear how to generalize the methodology to multi-source contexts, particularly when data sharing is restricted by privacy constraints.
1.1 Our contributions
In this paper, we propose two methods for multi-source causal survival analysis that accommodate events occurring in either continuous or discrete time. First, in settings where individual-level data can be shared and a CCOD assumption is adopted, we derive an estimator based on the semiparametric efficient influence function (EIF) (Section 2.2), which adjusts for covariate shift across sites while achieving both consistency and improved efficiency. Notably, our approach does not require the treatment or censoring mechanisms to be homogeneous across sources; nonetheless, we show that these nuisance functions can be estimated using pooled (global) models without sacrificing asymptotic guarantees. We establish uniform consistency and uniform asymptotic normality of the proposed estimator, thereby providing rigorous theoretical support for its use in multi-source survival settings.
Second, in more restrictive settings where individual-level data sharing is not allowed and the CCOD assumption may not hold, we introduce a federated learning framework (Section 2.3) that leverages a data-adaptive weighting scheme. This approach reweights source sites by up-weighting those more aligned with the target population and down-weighting—or even outright excluding—those likely to introduce bias, all without requiring access to individual-level data. Under a working partial CCOD assumption, we derive site-specific EIF-based estimators and show that all necessary quantities can be computed locally at each site, save for a density ratio term that can be consistently estimated using coarsened summary statistics. To aggregate the local estimators, we formulate and solve a constrained convex optimization problem to obtain federated weights that adaptively emphasizing informative sources while attenuating the influence of less relevant or biased sites. We establish the consistency and asymptotic normality of the proposed federated estimator and provide sufficient conditions under which it achieves efficiency gains over target-only estimation, i.e. its asymptotic variance is strictly smaller than that of an estimator relying solely on the target site.
For both approaches, we develop practical estimators that integrate cross-fitting [8] and off-the-shelf ensemble machine learning methods for nuisance function estimation [45, 17]. By leveraging flexible machine learning tools, we avoid reliance on restrictive parametric or semiparametric assumptions, while still achieving parametric convergence rates under standard regularity conditions.
We assess the empirical performance of our methods through comprehensive simulation studies and an application to the Antibody Mediated Protection (AMP) trials–a set of parallel, multi-site randomized controlled trials conducted across the Americas, Europe, and sub-Saharan Africa to evaluate the efficacy of the broadly neutralizing antibody VRC01 in preventing HIV-1 acquisition [12]. Our results demonstrate that the proposed estimators enable valid, efficient, and privacy-preserving causal effect estimation in complex multi-source survival settings, even under distributional shifts.
@@ 2.1 Notation, estimand, and assumptions
Consider $K$ studies, each of which could be randomized or observational. For each participant in a given study, we observe a covariate vector $\mathbf{X}$ representing baseline characteristics, and a binary treatment variable $A\in\{0,1\}$, where 0 denotes control and 1 denotes the active treatment. Under the Neyman-Rubin potential outcomes framework [38, 42], we assume each participant has two potential event times, $T^{(1)}$ and $T^{(0)}$, and two potential censoring times, $C^{(1)}$ and $C^{(0)}$, under treatments $A=1$ and $A=0$, respectively. Under the stable-unit treatment value assumption (SUTVA) [41], we have $T=T^{(A)}=AT^{(1)}+(1-A)T^{(0)}$ and $C=C^{(A)}=AC^{(1)}+(1-A)C^{(0)}$ for each participant; under right-censoring, we observe only the possibly censored outcome $Y=\min(T,C)$, together with an event indicator $\Delta=\mathbb{I}(T\leq C)$, where $\mathbb{I}(\cdot)$ is the indicator function. We denote an arbitrary observation as $\mathcal{O}$ and the observed data of sample size $n$ from all study sites as $\{\mathcal{O}_{i}=(\mathbf{X}_{i},A_{i},Y_{i},\Delta_{i},R_{i}),i=1,\dots,n\}$, where $R$ is a site variable taking values in $\{0,1,\dots,K-1\}$, with 0 representing the target site and the remaining $K-1$ values representing source sites, $K>1$.
In this paper, our estimand of interest is the target–site, treatment–specific survival function:
$\theta^{0}(t,a)=\mathbb{P}(T^{(a)}>t\mid R=0),\quad a\in\{0,1\},\quad t\in[0,\tau],$
for some positive $\tau<\infty$. Here, the upper subscript 0 of $\theta^{0}(t,a)$ indicates the target site $R=0$, and $\mathbb{P}$ denotes the distribution of an ideal data unit.
@@
is the conditional cumulative hazard function under $R=k$, with $N^{k}_{\delta}(t\mid a,\mathbf{x})=\mathbb{P}(Y\leq t,\Delta=\delta\mid A=a,\mathbf{X}=\mathbf{x},R=k)$ the conditional cumulative incidence function of event $(\delta=1)$ or censoring $(\delta=0)$, and $D^{k}(t\mid a,\mathbf{x})=\mathbb{P}(Y\geq t\mid A=a,\mathbf{X}=\mathbf{x},R=k)$. Under discrete time the product integral reduces to an ordinary product $\prod$, and under continuous time to $\exp\{-\Lambda^{k}(t\mid a,\mathbf{x})\}$.
Throughout, we make the following assumptions.
@@
Assumption 2.2 (Treatment-specific non-informative censoring).
@@
Assumption 2.3 (Positivity).
There exists an $\eta\in(0,\infty)$, such that $\mathbb{P}(R=k)\geq 1/\eta$ and for $\mathbb{P}$-almost all $\mathbf{x}$, $\min\cup_{k=0}^{K-1}\{\pi^{k}(a\mid\mathbf{x}),G^{k}(t\mid a,\mathbf{x})\}\geq 1/\eta$ and $\min_{k}S^{k}(t\mid a,\mathbf{x})>0$. Here, $\pi^{k}(a\mid\mathbf{x})=\mathbb{P}(A=a\mid\mathbf{X}=\mathbf{x},R=k)$ is the propensity score of the treatment $A=a$ and $G^{k}(t\mid a,\mathbf{x})=\mathbb{P}(C>t\mid A=a,\mathbf{X}=\mathbf{x},R=k)$ is the conditional survival function of the censoring time in site $k$, for $k=0,1,\dots,K-1$, $a\in\{0,1\}$ and $t\in[0,\tau]$.
The condition $\mathbb{P}(R=k)\geq 1/\eta$ in Assumption 2.3 ensures that the sampling or survey process prevents any site from contributing only an extremely small proportion of participants [32, 7].
Throughout, we define the following notation $\mathcal{H}_{t,a}$ for convenience:
$\mathcal{H}_{t,a}(\mathcal{O};S,G)=\frac{\mathbb{I}(Y\leq t,\delta=1)}{S(Y\mid a,\mathbf{X})G(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda(du\mid a,\mathbf{X})}{S(u\mid a,\mathbf{X})G(u\mid a,\mathbf{X})}.$ (1)
Since $S$ and $\Lambda$ have a one-to-one mapping, we did not involve $\Lambda$ in $\mathcal{H}_{t,a}(\mathcal{O};S,G)$. If we only use data from the target site $(R=0)$, the EIF of $\theta^{0}(t,a)$ can be given in Proposition 2.4.
Proposition 2.4 (Westling et al. [48]).
The nonparametric EIF for $\theta^{0}(t,a)$ given $t\in[0,\tau]$ and $a\in\{0,1\}$ is $\varphi^{*0}_{t,a}(\mathcal{O};S^{0},G^{0},\pi^{0})=\varphi^{*0}_{t,a}(\mathcal{O};\mathbb{P})=\varphi^{0}_{t,a}(\mathcal{O};\mathbb{P})-\theta^{0}(t,a)$, where
$\varphi^{0}_{t,a}(\mathcal{O};\mathbb{P})=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\bigg{[}1-\frac{\mathbb{I}(A=a)}{\pi^{0}(a\mid\mathbf{X})}\mathcal{H}_{t,a}(\mathcal{O};S^{0},G^{0})\bigg{]}S^{0}(t\mid a,\mathbf{X}).$
Throughout, let $\mathbb{P}_{n}[f(\mathcal{O})]=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ be the empirical average and $\widehat{\mathbb{P}}$ denote that the nuisance functions are replaced by their estimates. The above EIF motivates a doubly robust estimator $\widehat{\theta}^{0}_{n}(t,a)=\mathbb{P}_{n}[\widehat{\varphi}^{0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})]$. Under regularity conditions, $\widehat{\theta}^{0}_{n}(t,a)$ is regular and asymptotically linear (RAL) and attains the nonparametric efficiency bound uniformly over $t\in[0,\tau]$.
However, that result uses only data from the target site $R=0$. To further improve precision by borrowing strength from the $K-1$ source sites, we develop two strategies in the next two sections.
@@ 2.2 Efficient estimation under CCOD and data sharing
When target and source sites are sufficiently similar, it is often plausible to posit a common conditional outcome distribution (CCOD).
@@
Assumption 2.5 implies $S^{k}(t\mid a,\mathbf{X})=\bar{S}(t\mid a,\mathbf{X})$ for all sites $k$, while still allowing arbitrary shifts in the covariate distribution. Under this CCOD assumption, data from the source sites can be incorporated to yield more efficient estimators. We begin by extending Proposition 2.4 to this setting. Figure 1 presents the directed acyclic graph (DAG) for our full data structure, highlighting the roles of covariates, treatment, site, outcome, and censoring under CCOD.
Figure 1: DAG for data structures under different CCOD assumptions. When CCOD holds (Panel (a)), $R$ and $T$ are conditionally independent given treatment $A$ and covariates $\mathbf{X}$, consistent with Assumption 2.5. When CCOD is potentially violated, as highlighted by the orange dashed arrow in Panel (b), $R$ and $T$ are not necessarily conditionally independent.
Consider the following “global” nuisance functions (each defined without conditioning on $R$): the global propensity score $\bar{\pi}(a\mid\mathbf{x})=\mathbb{P}(A=a\mid\mathbf{X}=\mathbf{x})$, the global conditional survival function for the event time $\bar{S}(t\mid a,\mathbf{x})=\mathbb{P}(T>t\mid A=a,\mathbf{X}=\mathbf{x})$, and the global conditional survival function for the censoring time $\bar{G}(t\mid a,\mathbf{x})=\mathbb{P}(C>t\mid A=a,\mathbf{X}=\mathbf{x})$. Moreover, let the target site propensity be given by $q^{0}(\mathbf{X})=\mathbb{P}(R=0\mid\mathbf{X})$. Similar to Assumption 2.3, we impose the following positivity assumptions for these nuisance functions.
@@
There exists $\eta\in(0,\infty)$ such that for $\mathbb{P}$-almost all $\mathbf{x}$, $\min\{\bar{\pi}(a\mid\mathbf{x}),\bar{G}(t\mid a,\mathbf{x})\}\geq 1/\eta$ and $\bar{S}(t\mid a,\mathbf{x})>0$, for $a\in\{0,1\}$ and $t\in[0,\tau]$.
@@
Under Assumption 2.5, the semiparametric EIF of $\theta^{0}(t,a)$ given $t\in[0,\tau]$ and $a\in\{0,1\}$ is $\varphi^{*\text{CCOD}}_{t,a}(\mathcal{O};\bar{S},\bar{G},\bar{\pi})=\varphi^{*\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})=\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})-\theta^{0}(t,a)$, where
$\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})=\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}-\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\mathcal{H}_{t,a}(\mathcal{O};\bar{S},\bar{G})\bigg{]}\bar{S}(t\mid a,\mathbf{X}).$
@@
We employ $M$-fold cross-fitting following the DML framework [8], using flexible ensemble learning via the survSuperLearner [48] and SuperLearner [45] packages to estimate the nuisance functions. In Theorem 2.8, we demonstrate the (uniform) RAL property of the proposed CCOD estimator.
Theorem 2.8.
If Conditions D.1D.3 in Appendix D.1 (i.e., consistency, positivity, and rate condition on the product of prediction errors for the nuisance functions) hold, then $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{t,a})+o_{p}(n^{-1/2})$. In particular, $n^{1/2}(\widehat{\theta}^{\text{CCOD}}_{n}(t,a)-\theta^{0}(t,a))$ convergences in distribution to a normal random variable with mean zero and variance $\sigma^{2}=\mathbb{P}[(\varphi^{*\text{CCOD}}_{t,a})^{2}]$. If Conditions D.4D.5 (which strengthen the consistency and product error conditions to hold uniformly) also hold, then